OpenStudy (anonymous):

The formula for the maximum focus range or hyperfocal distance, H, of a camera is given by H= L^2/F(C) , where L is the focal length of the lens, f is the setting on the camera to reduce or enlarge the aperture, and c is the diameter of the circle of least confusion. The formula can be rewritten to solve for L as L=F(C)/H^2 TRUE FALSE

OpenStudy (jack1):

H= L^2/F(C) solve for L \(\Large H= \frac {L^2}{F(C)}\) \(\large H \times F(C) = L^2\) \(\large L^2 = H \times F(C) \) \(\large L = \sqrt {H \times F(C)} \) i think...?